OpenStudy (anonymous):

if f(x)=lx-4l,which of the following has the least value? a)f(8) b)f(4) c)f(0) d)f(-4) e)f(-8)

OpenStudy (mayankdevnani):

option e

OpenStudy (mayankdevnani):

@umar284

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

You can plug the values and compare the answers: For example: f(x)=lx-4l f(8) = |8-4| = |4| = 4 f(-8) = |-12 -4| = |-16| = 16 So, you can see that f(-8) is definitely NOT the answer! Can you do it now?

OpenStudy (anonymous):

im confused by the modulus

OpenStudy (anonymous):

@RolyPoly

OpenStudy (anonymous):

I'm sorry! I wasn't here! Which part don't you understand??

OpenStudy (mayankdevnani):

oh!! sry its absolute value